OpenStudy (alfordabby):

A delivery truck is transporting boxes of two sizes: large and small. The large boxes weigh 45 pounds each, and the small boxes weigh 25 pounds each. There are 115 boxes in all. If the truck is carrying a total of 3975 pounds in boxes, how many of each type of box is it carrying?

OpenStudy (alfordabby):

x + y = 115 ? 45x + 25y = 2975 ?

OpenStudy (alfordabby):

is my equation right?

OpenStudy (alfordabby):

45x + 25(115 - x) = 3975 45x + 2875 - 25x = 3975 20x + 2875 = 3975 20x = 1100 x = 55

OpenStudy (alfordabby):

x + y = 115 55 + y = 115 y = 60